Ms. Brown, the owner of Southern Comfort Inc., has approached you with an offer to sell you her company for paid her a dividend of $10 ilion last year. She provided you with the following expected cash flow analysis (see the table below), and your assistant confirmed that the projections are reasonable. T3 $17 $68 Year TO T1 T2 ValueS12 If Ms Brown sells to you, what is her realized annual rate of return on her original A. investment? The risk free rate is 5%. INI). If you negotiate and are able to buy Southern Comfort for $75 million, what is your expected rate of return on this investment? Should you buy Southern Comfort? (answer YES or C.
Answer #1
A Original Investment $60 million
B Dividend received in year 1 $10 million
C Sales value in year 2 $80 million
D=B+C Total amount received $90 million
E=D-A Realized return in two years $30 million
F=E/2 Average annual return $15 million
G=F/A Realized annual rate of return 0.25
Realized annual rate of return 25.00%
Present Value (PV) of Cash Flow:
(Cash Flow)/((1+i)^N)
i=Discount Rate=Risk free rate=5% 0.05
N=Year of Cash Flow
N Year 1 2 3
A Dividend $12 $15 $17
B Firm value $68
C=A+B Total cash in flow($million) $12 $15 $85 SUM
D=C/(1.05^N) Present Value of Cash in flow 11.42857 13.60544 73.4262 98.46021
Present Value of Cash in flow $98.46 million
Net Present Value (NPV) $18.46 million (98.46-80)
NPV is positive
YES, We should buy for $80 million
If we buy for 75 million,
Total realized amount in 3 years $112 million (12+15+17+68)
Total realized return in 3 years $37 million (112-75)
Average annual return $12.33 million (37/3)
Annual rate of return           0.16 (12.33/75)
Annual rate of return 16.44%